OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

OpenStudy (anonymous):

I think I'm supposed to solve this in the form of a rational function.

OpenStudy (anonymous):

I'm just not sure how I'm supposed to form my response. I originally did: 1/(x-8) + 1/x = 1/3 But, I wasn't sure how to simplify and I'm not sure if I even did it right.

OpenStudy (anonymous):

if we recall: harry x/8 hours jimmy x hours x + x/8 = 3 hours \[\frac{ 9x }{ 8 }=3\] 9x = 24 x = 24 /9 jimmy takes 2,66 hours